Hi, I need help with this problem: Your older sister is getting married and to save money she is making her own cylinder shaped cake with three layers. Top layer: Radius= 3cm Height = 3 cm Middle Layer: Radius= 6cm Height = 4 cm Bottom layer: Radius = 10cm Height= 5cm If each tub of frosting costs $6.99 and a each tub of frosting covers 120 cm^2 of cake, figure out how much money your sister will have to spend on frosting.
Answer by nerdybill(7384)

Your older sister is getting married and to save money she is making her own cylinder shaped cake with three layers. Top layer: Radius= 3cm Height = 3 cm Middle Layer: Radius= 6cm Height = 4 cm Bottom layer: Radius = 10cm Height= 5cm If each tub of frosting costs $6.99 and a each tub of frosting covers 120 cm^2 of cake, figure out how much money your sister will have to spend on frosting
.
Surface Area = Area of top +Area of the side
.
Surface Area = (Area of top) + (perimeter of top)* height
.
Surface Area = (pi r^2) + (2 pi r)* h
.
Top layer:
Surface Area = (3.14*3^2) + (2*3.14*3)* 3
Surface Area = (3.14*9) + (6*3.14)* 3
Surface Area = 84.78 cm^2
.
Middle layer:
Surface Area = (3.14*6^2) + (2*3.14*6)* 4
Surface Area = (3.14*36) + (12*3.14)* 4
Surface Area = 263.76 cm^2
.
Bottom layer:
Surface Area = (3.14*10^2) + (2*3.14*10)* 5
Surface Area = (3.14*100) + (20*3.14)* 5
Surface Area = 628 cm^2
.
Total area = 976.54 cm^2
.
Deciding how many tubs:
976.54/120 = 8.14 tubs
.
Since you can't buy a partial tub, you must buy the next whole number up:
9 tubs of frosting
.
Total cost:
9 * 6.99 = $62.91